Holtzman Clothiers's stock currently sells for $15.00 a share. It just paid a dividend of $3.00 a share (i.e., D0 = $3.00). The dividend is expected to grow at a constant rate of 8% a year.

What stock price is expected 1 year from now? Round your answer to the nearest cent. ?
$   

What is the required rate of return? Do not round intermediate calculations. Round your answer to two decimal places. ?
Answer #1

This question requires application of constant growth dividend discount model according to which: Div Po r-g Po Price of Stoc

Div1 = $3 * (1 + 8%) = $3.24

15 = \frac{3.24}{r - 0.08}

r - 0.08 = 0.216

r = 29.60% ---> Required rate of Return

P_1 = \frac{Div_2}{r - g}

Div2 = $3.24 * (1 + 8%) = $3.4992

P_1 = \frac{3.4992}{0.216 - 0.08}

P_1 = 25.73 ----> Price of stock in 1 year